Prove that sin60/cos60=tan60. Please explain every step
Answer by Alan3354(30993)

Prove that sin60/cos60=tan60. Please explain every step
-----------------
Use the unit circle.
sin = y/r = y
cos = x/r = x
---
sin/cos = y/x = tan (of any angle)